OpenStudy (anonymous):

Evaluate. –317 – 221 – (–354)

Mehek (mehek14):

– (–354) means add 354 so -317 - 221 + 354 what do you get?

OpenStudy (anonymous):

wait am i doing -317 MINUS 221 + 354?

Mehek (mehek14):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1088?

Mehek (mehek14):

how did you get that???

OpenStudy (anonymous):

well 317 was negative so i had to add not minus then i added WAIT one sec

OpenStudy (anonymous):

-184

Mehek (mehek14):

-317-221 = -538 -538 + 354 = -184

Mehek (mehek14):

yes correct

OpenStudy (anonymous):

sorry i got mixed up lol

OpenStudy (anonymous):

well thank you!

Mehek (mehek14):

it's ok but make sure you don't get mixed up again and your welcome :)